OpenStudy (anonymous):

x+2y=-3 x-4y=15 Using the substitution method, what do x and y equal for the equations?

OpenStudy (anonymous):

x= -3-2y x= 15 + 4y -3-2y = 15 + 4y -6y = 18 y = -3 x= 15+ 4(-3) x= 3